OpenStudy (anonymous):

find two numbers whose sum is 28 and whose product is 96.

OpenStudy (anonymous):

let the 2 numbers be x and y. x+y = 28 xy = 96

OpenStudy (anonymous):

x = 28 - y So, y(28-y) = 96 28y-y^2 = 96 y^2 - 28y +96 = 0 (y-24)(y-4)=0 y=4,24 When y = 4, x = 24 When y = 24, x = 4 So the 2 numbers are 4 and 24.